How many moles and how many atoms of nickel are in one nickel


1
Expert's answer
2021-01-25T04:31:28-0500

How many moles and how many atoms of nickel are in one gram of nickel?


Solution:

1 mol Ni is equal to 58.6934 grams of Ni

Hence,

Moles of Ni = 1 g × (1 mol / 58.6934 g) = 0.017 moles

Moles of Ni = 0.017 moles

1 gram Ni equal to 0.017 Ni moles


One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

Number of Ni atoms = 0.017 mol × (6.022×1023 atoms / 1 mol) = 1.02×1022 Ni atoms

Number of Ni atoms = 1.02×1022 atoms


Answer: 1 gram of Ni equal to 0.017 moles of Ni, or 1.02×1022 Ni atoms.